Storrs Cycles has just started selling the new Cyclone mountain bike, with monthly sales as shown in the table. Month Sales Jplease answer part 2 of C
Answer #1

To calculate the mean absolute deviation We have to first calculate the deviation and absolute deviation for all the periods. Where,

Deviation = Actual value - forecasted value

Absolute deviation = absolute value of deviation

So using the above formula the deviation and absolute deviation for all the periods are

  • For April deviation = 365-398.00 = -33 and absolute deviation = 33
  • For May deviation = 400-386.33 = 13.67 and absolute deviation = 13.67

Mean absolute deviation = Sum of the absolute deviation for all the periods/number of periods

= (33+13.67)/2

= 46.67/2

= 23.34